# Crossings and diffusion in Poisson driven marked random connection models

### v1

@@ 5 Connections, seeds and sprinkling
where $G(A,\tilde{\mathbb{X}})$ denotes the bipartite RCM with connection function $\varphi$, i.e. the graph with vertex set $A\cup\tilde{\mathbb{X}}$ where an edge $\{t,t^{\prime}\}$ is created with probability $\varphi(t,t^{\prime})$ independently when varying $t\in A$ and $t^{\prime}\in\tilde{\mathbb{X}}$.
@@
In what follows, to shorten, we will write PPP$[\mathfrak{m}]$ for a PPP with intensity measure $\mathfrak{m}$. If $A\cap B\not=\varnothing$, then (17) is trivially true. Let us then restrict to $A\cap B=\varnothing$. Given a realization of ${\mathbb{X}}$, a point $t\in{\mathbb{X}}$ is adjacent to $A$ in $G({\mathbb{X}}\cup A)$ with probability $1-\bar{\varphi}(t,A)$ (the probability is w.r.t. the connection random variables) and this takes place independently for each point $t\in{\mathbb{X}}$. Hence, if we set $\xi:=\{t\in{\mathbb{X}}\,:\,t\not\sim A\text{ in }G({\mathbb{X}}\cup A)\}$ and $\eta:=\{t\in{\mathbb{X}}\,:\,t\sim A\text{ in }G({\mathbb{X}}\cup A)\}$, we get that $\xi$ and $\eta$ are independent, $\xi\sim\text{PPP}[\rho\bar{\varphi}(t,A)\mu(\mathrm{d}t)]$ and $\eta\sim\text{PPP}[n(\mathrm{d}t)]$, where $n(\mathrm{d}t):=\rho(1-\bar{\varphi}(t,A))\mu(\mathrm{d}t)$. Since, in (17), we may assume that ${\mathbb{X}}$ and $\tilde{\mathbb{X}}$ are independent, we may set ${\mathbb{Y}}=\xi$ and $G({\mathbb{Y}})$ to be the restriction of $G({\mathbb{X}})$ to the vertex set $\xi$.
@@
${\mathbb{P}}\left(A\leftrightarrow W\cup B\text{ in }G\big{(}A,\tilde{\mathbb{X}}\big{)}\cup G\big{(}\tilde{\mathbb{X}}\cup W\big{)}\right)\\ \geq{\mathbb{P}}\left(\exists t\in\tilde{\mathbb{X}}:t\sim A\text{ in }G(\{t\}\cup A)\text{ and }\left(t\in B\text{ or }t\sim W\text{ in }G(\{t\}\cup W)\right)\right)=1-{\mathbb{E}}\left[e^{-\delta Z(W)}\right].$ (20)
@@
$\Big{\{}t\in\tilde{\mathbb{X}}\cap B:t\sim A\text{ in }G(\{t\}\cup A)\Big{\}},\quad\Big{\{}t\in\tilde{\mathbb{X}}\setminus B:t\sim A\text{ in }G(\{t\}\cup A)\text{ and }t\sim W\text{ in }G(\{t\}\cup W)\Big{\}}\,,$
@@
Fix $r<s-1$ and $x\in{\mathbb{R}}^{d}$. Let $G=(\omega,E)$ be a random graph with $\omega\in\Omega$. Let $G^{\prime}$ be the restriction of $G$ to the vertex set $\omega\cap(\Lambda_{s}(x)\times{\mathbb{M}})$. We say that $\mathcal{A}(r,s,x)$ occurs for $G$, if, for any two paths $\gamma,\gamma^{\prime}$ in $G^{\prime}$ from $\Lambda_{r}(x)\times{\mathbb{M}}$ to $\partial_{*}\Lambda_{s}(x)\times{\mathbb{M}}$, there is a path in $G^{\prime}$ from a vertex of $\gamma$ to a vertex of $\gamma^{\prime}$. We set $\mathcal{A}(r,s):=\mathcal{A}(r,s,0)$.
@@
As in [DembinNaN]*Lemma 6.3, for $r<s-1$ we define $\mathcal{E}_{r,s}$ as the event that the graph $G({\mathbb{X}})$ restricted to $\Lambda_{s}\times{\mathbb{M}}$ has at least two connected components intersecting both $\Lambda_{r}\times{\mathbb{M}}$ and $\partial_{*}\Lambda_{s}\times{\mathbb{M}}$. Note that $\mathcal{A}(r,s)\subset\mathcal{E}_{r,s}^{c}$. Moreover, if the event $\bigcap_{s=r+1}^{\infty}\mathcal{E}_{r,s}$ occurs, then the finite family of the clusters of vertices in $\Lambda_{r}\times{\mathbb{M}}$ must contain at least two disjoint infinite clusters. By Assumption 2.3(iv) we then get that $\lim_{r\to+\infty}\lim_{s\to+\infty}{\mathbb{P}}(\mathcal{E}_{r,s})={\mathbb{P}}\left(\bigcup_{r=1}^{\infty}\big{(}\bigcap_{s=r+1}^{\infty}\mathcal{E}_{r,s}\big{)}\right)=0$ and therefore $\lim_{r\to+\infty}\lim_{s\to+\infty}{\mathbb{P}}(\mathcal{A}(r,s))=1$. This easily allows to take $r_{0}\leq k\ll K\ll n\ll N$ satisfying (25), (26) and (27). ∎
@@
$\mathcal{E}_{i}$ :=${}:=\left\{(x_{i},m_{i})\leftrightarrow\partial_{*}Q_{i}\times{\mathbb{M}}\text{ in }G({\mathbb{X}}\cup S)\right\}\cap\mathcal{A}(k,K,x_{i})\,,$
$\mathcal{B}_{i}$ :=${}:=\left\{Q_{i}^{\prime}\times{\mathbb{M}}\not\leftrightarrow\partial_{*}\Lambda_{N}\times{\mathbb{M}}\text{ in }G({\mathbb{X}})\right\}\,,$
@@
To estimate ${\mathbb{P}}(\mathcal{B}_{i})$ one cannot apply (27) directly, because $Q_{i}^{\prime}$ is not necessarily centered at 00. As in [Duminil-Copin21], we fix $j(i)\in\{1,\dots,d\}$ and $\sigma(i)\in\{\pm 1\}^{d}$, so that $F^{i}=x_{i}+F^{j(i),\sigma(i)}_{N}$ (recall (16)) does not intersect the interior of $\Lambda_{N-1}$. Note that the event $\mathcal{D}_{i}:=\{Q_{i}^{\prime}\times{\mathbb{M}}\leftrightarrow F^{i}\times{\mathbb{M}}\text{ in }\Lambda_{N}(x_{i})\times{\mathbb{M}}\text{ in }G({\mathbb{X}})\}$ is disjoint from $\mathcal{B}_{i}$. Hence,
@@
$\{S\leftrightarrow\partial_{*}\Lambda_{N}\times{\mathbb{M}}\text{ in }G({\mathbb{X}}\cup S)\}\,,\qquad\{\Lambda_{k}\times{\mathbb{M}}\leftrightarrow F_{N}\times{\mathbb{M}}\text{ in }\Lambda_{N}\times{\mathbb{M}}\text{ in }G({\mathbb{X}})\}\,,\qquad\mathcal{A}(n,N),$
@@
For the next result recall that $\lambda^{\prime}$ is a fixed density with $\lambda^{\prime}>\lambda$ (see Section 4.4).
@@
    the event $\mathcal{H}$ is of the form :=$\mathcal{H}:=\{(\omega,\underline{u})\in\Omega\times[0,1]^{J}:G(\omega,\underline{u})\in\mathcal{H}_{*}\}$ for a suitable set $\mathcal{H}_{*}$ of graphs with vertex set in ${\mathbb{R}}^{d}\times{\mathbb{M}}$. $\mathcal{H}_{*}$ is decreasing w.r.t. inclusion (i.e. if $\mathcal{G}\in\mathcal{H}_{*}$ and $\mathcal{G}\supset\mathcal{G}^{\prime}$, then $\mathcal{G}^{\prime}\in\mathcal{H}_{*}$) and a graph on ${\mathbb{R}}^{d}\times{\mathbb{M}}$ belongs to $\mathcal{H}_{*}$ if and only if its restriction to $\Lambda_{6N}\times{\mathbb{M}}$ belongs to $\mathcal{H}_{*}$.
@@
(the above identity follows from Mecke equation, see [Last18]*Theorem 4.1). By Lemma 5.3, we have $12^{d}\rho\leq 2^{d}\lambda^{\prime}c<$ 1 (by choosing $c$ small), so (32) can be made arbitrarily small by taking $n$ large.
@@
Let $\mathcal{H}$ be the event such that its complement $\mathcal{H}^{c}$ is the following: there are at least $3]{N}$ distinct bad boxes $\Lambda_{n}(y_{1}),\Lambda_{n}(y_{2}),\dots,\Lambda_{n}(y_{\sqrt[3]{N}})$ with successive centers in $n{\mathbb{Z}}^{d}$ at $\ell^{\infty}$-distance at most $4n$ (to simplify the notation we assume $3]{N}$ to be integer). The event $\mathcal{H}$ clearly satisfies (iii) as $\mathcal{H}^{c}$ is described in terms of $G(\omega,\underline{u})$ and is increasing w.r.t. graph inclusion.
@@
Let us now prove (i). Since good boxes form a 1-dependent process, the Liggett–Schonmann–Stacey theorem [Liggett99] (also see [Grimmett99]*Theorem (7.65)) implies the following. For $p\in(0,1)$, by taking the rightmost term in (32) small enough (that is, $n$ large enough), the random field $(T_{y})_{y\in{\mathbb{Z}}^{d}}$ with $T_{y}:=\mathbbm{1}(\Lambda_{n}(ny)\text{ is good})$ stochastically dominates a Bernoulli site percolation with parameter $p$. Due to the this stochastic domination and since bad boxes have to intersect $\Lambda_{6N}$, we have $P(\mathcal{H})\geq 1-\varepsilon$, taking $p$ close to 1 (and therefore $n$ large) and afterwards $N$ large enough. This follows easily since for the above Bernoulli site percolation the probability to have a path of length $3]{N}$ of closed points in $\Lambda_{\lceil 6N/n\rceil+1}$ such that consecutive points have $\ell^{\infty}$–distance at most $4$ is upper bounded by $(12N+5)^{d}\left(9^{d}(1-p)\right)^{\sqrt[3]{N}}$.
@@
We next prove (ii). We first observe that the Voronoi tessellation associated to $n{\mathbb{Z}}^{d}$ is given by the boxes $\Lambda_{n/2}(z)$ with $z\in n{\mathbb{Z}}^{d}$. In particular, for each $x\in{\mathbb{R}}^{d}$ we can find $z\in n{\mathbb{Z}}^{d}$ such that $x\in\Lambda_{n/2}(z)$. Fix $(\omega,\underline{u})\in\mathcal{H}$ such that there exists a path $\gamma=(x_{i},m_{i})_{i=0}^{r}$ as in the statement of (ii). Let $z_{i}\in n{\mathbb{Z}}^{d}$ be such that $x_{i}\in\Lambda_{n/2}(z_{i})$. Recall that $|x_{i+1}-x_{i}|_{\infty}\leq 1$ for any consecutive points in $\gamma$ (as ${\ell}_{*}=1$) and $|x_{r}-x_{0}|_{\infty}\geq\sqrt{N}-1$. Then we prune the finite sequence $\Lambda_{n}(z_{1}),\Lambda_{n}(z_{2}),\dots,\Lambda_{n}(z_{r})$, by loop-erasing and by removing intersecting boxes, and we extract distinct points $y_{0},y_{1},\dots,y_{s}$ among $z_{0},z_{1},\dots,z_{r}$ such that $|y_{j+1}-y_{j}|_{\infty}=2n$ for each $j=0,1,\dots,s-1$, where $y_{0}:=z_{0}$, and $|y_{s}-z_{r}|_{\infty}\leq 2n$. Note that necessarily $s\geq(\sqrt{N}-1-2n)/(2n)$. For each $j=1,2,\dots,s$ we note that the path $\gamma$ goes from $(x_{0},m_{0})$ into $\Lambda_{n/2}(y_{j})\times{\mathbb{M}}$, hence $\gamma$ needs to give rise to a path $\gamma_{j}$ in $\Lambda_{n}(y_{j})\times{\mathbb{M}}$ of $\ell^{\infty}$-diameter at least $n/2-1$. We claim that some $\gamma_{j}$ contains a seed.
@@ 6.1 Order $\prec$ on $\Delta\{x_{1},\dots,x_{n}\}$
The above order is the one introduced in [Tanemura93]*Section 4 with a relevant exception. Indeed in [Tanemura93] the author supposes that $(x_{1},x_{2},\dots,x_{n})$ is a path in ${\mathbb{Z}}^{2}$ but then refers to the ordering also when dealing with more general strings, as the above ones. If $(x_{1},x_{2},\dots,x_{n})$ is a path, then $a(k)=k-1$ for all $k=1,2,\dots,n$ and the above definition of $\prec$ coincides with the order on $\Delta\{x_{1},x_{2},\dots,x_{n}\}$ introduced in [Tanemura93]*Section 4.
@@ 6.2 Exploration process
We denote $x_{1}^{s}=(0,s)$ for $s\in\{0,\dots,M-1\}$. Note that in the above notation $\Lambda_{L}^{\prime}$, $\partial_{l}\Lambda_{L}^{\prime}$ and $\partial_{r}\Lambda_{L}^{\prime}$ the integer $M$ is kept implicit. Suppose that on some probability space, with probability measure ${\mathbb{Q}}_{L}$ the following events are defined: $\{x_{1}^{s}:=(0,s)\text{ is occupied}\}$ for $s=0,1,\dots,M-1$, $\{y\text{ is linked to }x\}$ for $|x-y|=1$ and $x,y\in\Lambda^{\prime}_{L}$.
@@
$\begin{cases}\text{if $x^{s}_{1}$ is occupied, then $C^{s}_{1}=\left(E^{s}_{1},F^{s}_{1}\right):=(\{x^{s}_{1}\},\emptyset)$}\,,\\ \text{if $x^{s}_{1}$ is not occupied, then $C^{s}_{1}=\left(E^{s}_{1},F^{s}_{1}\right):=(\emptyset,\{x^{s}_{1}\})$}\,.\end{cases}$ (35)
@@
If $x_{1}^{s}$ is occupied (i.e. $E^{s}_{1}\not=\emptyset$), then we proceed as follows. Suppose that we have defined all the sets preceding $C^{s}_{j+1}$ in the above string (34) (i.e. up to $C^{s}_{j}$), that we have not assigned to $J^{s}$ any value yet and that we want to define $C^{s}_{j+1}$. We call $W^{s}_{j}$ the points of $\Lambda_{L}^{\prime}$ involved in the construction up to this moment, i.e.
@@
We recall that it must be $E^{s}_{0}\subset E^{s}_{1}\subset\cdots\subset E^{s}_{j}$ and at each inclusion either the two sets are equal or the second one is obtained from the first one by adding exactly one point adjacent to the previous one. Hence, we can order the elements of $E^{s}_{j}$ according to the order with which the points have been added. This order on $E^{s}_{j}$ induces an order $\prec$ on the boundary $\Delta E^{s}_{j}$ as described in Section 6.1. We call $\mathcal{P}^{s}_{j}$ the following property: $E^{s}_{j}$ is disjoint from the right boundary $\partial_{r}\Lambda_{L}^{\prime}$ and $(\Lambda_{L}^{\prime}\cap\Delta{E}^{s}_{j})\setminus W^{s}_{j}\not=\emptyset$. If property $\mathcal{P}^{s}_{j}$ is satisfied, then we define $x^{s}_{j+1}$ as the largest element of $(\Lambda_{L}^{\prime}\cap\Delta{E}^{s}_{j})\setminus W^{s}_{j}$ w.r.t. $\prec$. Let $k$ be the largest integer such that $x_{k}^{s}\in E^{s}_{j}$ and $|x^{s}_{j+1}-x^{s}_{k}|=1$. We decide where to add $x^{s}_{j+1}$ as follows:
$\begin{cases}\text{if $x^{s}_{j+1}$ is linked to $x^{s}_{k}$, then $C^{s}_{j+1}:=\big{(}E^{s}_{j}\cup\{x^{s}_{j+1}\},F^{s}_{j}\big{)}$,}\\ \text{if $x^{s}_{j+1}$ is not linked to $x^{s}_{k}$, then $C^{s}_{j+1}:=\left(E^{s}_{j},F^{s}_{j}\cup\{x^{s}_{j+1}\}\right)$\,.}\end{cases}$ (36)

### v2

@@ 5 Connections, seeds and sprinkling
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Fix r<s1r<s-1 and $x\in{\mathbb{R}}^{d}$. Let $G=(\omega,E)$ be a random graph with $\omega\in\Omega$. Let $G^{\prime}$ be the restriction of $G$ to the vertex set $\omega\cap(\Lambda_{s}(x)\times{\mathbb{M}})$. We say that $\mathcal{A}(r,s,x)$ occurs for $G$, if, for any two paths $\gamma,\gamma^{\prime}$ in $G^{\prime}$ from $\Lambda_{r}(x)\times{\mathbb{M}}$ to $\partial_{*}\Lambda_{s}(x)\times{\mathbb{M}}$, there is a path in $G^{\prime}$ from a vertex of $\gamma$ to a vertex of $\gamma^{\prime}$. We set $\mathcal{A}(r,s):=\mathcal{A}(r,s,0)$.
We emphasise that we do not require and will not prove strong local uniqueness, namely e.g. $\liminf_{N\to\infty}P_{\rho,\nu}\otimes Q(\mathcal{A}(3N,4N))=1$. If this result were available a priori, it would greatly facilitate the proof, in particular essentially removing the need of Section 6, as discussed in Section 1.3. However, proving such strong local uniqueness often requires a significant amount of work even after Theorem 2.6 is proved. This is why we have opted for the simpler approach presented here.
@@
As in [DembinToappear]*Lemma 6.3, for r<s1r<s-1 we define $\mathcal{E}_{r,s}$ as the event that the graph $G({\mathbb{X}})$ restricted to $\Lambda_{s}\times{\mathbb{M}}$ has at least two connected components intersecting both $\Lambda_{r}\times{\mathbb{M}}$ and $\partial_{*}\Lambda_{s}\times{\mathbb{M}}$. Note that $\mathcal{A}(r,s)\subset\mathcal{E}_{r,s}^{c}$. Moreover, if the event $\bigcap_{s=r+1}^{\infty}\mathcal{E}_{r,s}$ occurs, then the finite family of the clusters of vertices in $\Lambda_{r}\times{\mathbb{M}}$ must contain at least two disjoint infinite clusters. By Assumption 2.3(iv) we then get that $\lim_{r\to+\infty}\lim_{s\to+\infty}{\mathbb{P}}(\mathcal{E}_{r,s})={\mathbb{P}}\left(\bigcup_{r=1}^{\infty}\big(\bigcap_{s=r+1}^{\infty}\mathcal{E}_{r,s}\big)\right)=0$ and therefore $\lim_{r\to+\infty}\lim_{s\to+\infty}{\mathbb{P}}(\mathcal{A}(r,s))=1$. This easily allows to take $r_{0}\leq k\ll K\ll n\ll N$ satisfying (25), (26) and (27). ∎
@@
$\mathcal{E}_{i}$ ${}:=\left\{(x_{i},m_{i})\leftrightarrow\partial_{*}Q_{i}\times{\mathbb{M}}\text{ in }G({\mathbb{X}}\cup S)\right\}\cap\mathcal{A}(k,K,x_{i})\,$,
$\mathcal{B}_{i}$ ${}:=\left\{Q_{i}^{\prime}\times{\mathbb{M}}\not\leftrightarrow\partial_{*}\Lambda_{N}\times{\mathbb{M}}\text{ in }G({\mathbb{X}})\right\}\,$,
@@
To estimate ${\mathbb{P}}(\mathcal{B}_{i})$ one cannot apply (27) directly, because $Q_{i}^{\prime}$ is not necessarily centered at 0. As in [Duminil-Copin21], we fix $j(i)\in\{1,\dots,d\}$ and $\sigma(i)\in\{\pm 1\}^{d}$, so that $F^{i}=x_{i}+F^{j(i),\sigma(i)}_{N}$ (recall (16)) does not intersect the interior of $\Lambda_{N-1}$. Note that the event $\mathcal{D}_{i}:=\{Q_{i}^{\prime}\times{\mathbb{M}}\leftrightarrow F^{i}\times{\mathbb{M}}\text{ in }\Lambda_{N}(x_{i})\times{\mathbb{M}}\text{ in }G({\mathbb{X}})\}$ is disjoint from $\mathcal{B}_{i}$. Hence,
@@
$\{S\leftrightarrow\partial_{*}\Lambda_{N}\times{\mathbb{M}}\text{ in }G({\mathbb{X}}\cup S)\}\,,\hskip 18.49988pt\{\Lambda_{k}\times{\mathbb{M}}\leftrightarrow F_{N}\times{\mathbb{M}}\text{ in }\Lambda_{N}\times{\mathbb{M}}\text{ in }G({\mathbb{X}})\}\,,\hskip 18.49988pt\mathcal{A}(n,N),$
@@
For the next result recall that $\lambda^{\prime}$ is a fixed density with $\lambda^{\prime}>\lambda$ (see Section 4.4). This lemma is our starting point for dealing with the presence of bad marks in the model.
@@
    the event $\mathcal{H}$ is of the form $\mathcal{H}:=\{(\omega,\underline{u})\in\Omega\times[0,1]^{J}:G(\omega,\underline{u})\in\mathcal{H}_{*}\}$ for a suitable set $\mathcal{H}_{*}$ of graphs with vertex set in ${\mathbb{R}}^{d}\times{\mathbb{M}}$. $\mathcal{H}_{*}$ is decreasing w.r.t. inclusion (i.e. if $\mathcal{G}\in\mathcal{H}_{*}$ and $\mathcal{G}\supset\mathcal{G}^{\prime}$, then $\mathcal{G}^{\prime}\in\mathcal{H}_{*}$) and a graph on ${\mathbb{R}}^{d}\times{\mathbb{M}}$ belongs to $\mathcal{H}_{*}$ if and only if its restriction to $\Lambda_{6N}\times{\mathbb{M}}$ belongs to $\mathcal{H}_{*}$.
@@
(the above identity follows from Mecke equation, see [Last18]*Theorem 4.1). By Lemma 5.3, we have $2^{d}\rho\leq 2^{d}\lambda^{\prime}c<1$ (by choosing $c$ small), so (32) can be made arbitrarily small by taking $n$ large.
@@
Let $\mathcal{H}$ be the event such that its complement $\mathcal{H}^{c}$ is the following: there are at least $\sqrt[3]{N}$ distinct bad boxes $\Lambda_{n}(y_{1}),\Lambda_{n}(y_{2}),\dots,\Lambda_{n}(y_{\sqrt[3]{N}})$ with successive centers in $n{\mathbb{Z}}^{d}$ at $\ell^{\infty}$-distance at most $4n$ (to simplify the notation we assume $\sqrt[3]{N}$ to be integer). The event $\mathcal{H}$ clearly satisfies (iii) as $\mathcal{H}^{c}$ is described in terms of $G(\omega,\underline{u})$ and is increasing w.r.t. graph inclusion.
@@
Let us now prove (i). Since good boxes form a 1-dependent process, the Liggett–Schonmann–Stacey theorem [Liggett97] (also see [Grimmett99]*Theorem (7.65)) implies the following. For $p\in(0,1)$, by taking the rightmost term in (32) small enough (that is, $n$ large enough), the random field $(T_{y})_{y\in{\mathbb{Z}}^{d}}$ with $T_{y}:=\mathbbm{1}(\Lambda_{n}(ny)\text{ is good})$ stochastically dominates a Bernoulli site percolation with parameter $p$. Due to the this stochastic domination and since bad boxes have to intersect $\Lambda_{6N}$, we have $P(\mathcal{H})\geq 1-\varepsilon$, taking $p$ close to 1 (and therefore $n$ large) and afterwards $N$ large enough. This follows easily since for the above Bernoulli site percolation the probability to have a path of length $\sqrt[3]{N}$ of closed points in $\Lambda_{\lceil 6N/n\rceil+1}$ such that consecutive points have $\ell^{\infty}$–distance at most $4$ is upper bounded by $(12N+5)^{d}\left(9^{d}(1-p)\right)^{\sqrt[3]{N}}$.
@@
We next prove (ii). We first observe that the Voronoi tessellation associated to $n{\mathbb{Z}}^{d}$ is given by the boxes $\Lambda_{n/2}(z)$ with $z\in n{\mathbb{Z}}^{d}$. In particular, for each $x\in{\mathbb{R}}^{d}$ we can find $z\in n{\mathbb{Z}}^{d}$ such that $x\in\Lambda_{n/2}(z)$. Fix $(\omega,\underline{u})\in\mathcal{H}$ such that there exists a path $\gamma=(x_{i},m_{i})_{i=0}^{r}$ as in the statement of (ii). Let $z_{i}\in n{\mathbb{Z}}^{d}$ be such that $x_{i}\in\Lambda_{n/2}(z_{i})$. Recall that $|x_{i+1}-x_{i}|_{\infty}\leq 1$ for any consecutive points in $\gamma$ (as ${\ell}_{*}=1$) and $|x_{r}-x_{0}|_{\infty}\geq\sqrt{N}-1$. Then we prune the finite sequence $\Lambda_{n}(z_{1}),\Lambda_{n}(z_{2}),\dots,\Lambda_{n}(z_{r})$, by loop-erasing and by removing intersecting boxes, and we extract distinct points $y_{0},y_{1},\dots,y_{s}$ among $z_{0},z_{1},\dots,z_{r}$ such that $|y_{j+1}-y_{j}|_{\infty}=2n$ for each $j=0,1,\dots,s-1$, where $y_{0}:=z_{0}$, and $|y_{s}-z_{r}|_{\infty}\leq 2n$. Note that necessarily $s\geq(\sqrt{N}-1-2n)/(2n)$. For each $j=1,2,\dots,s$ we note that the path $\gamma$ goes from $(x_{0},m_{0})$ into $\Lambda_{n/2}(y_{j})\times{\mathbb{M}}$, hence $\gamma$ needs to give rise to a path $\gamma_{j}$ in $\Lambda_{n}(y_{j})\times{\mathbb{M}}$ of $\ell^{\infty}$-diameter at least n/21n/2-1. We claim that some $\gamma_{j}$ contains a seed.
@@ 6.1 Order $\prec$ on $\Delta\{x_{1},\dots,x_{n}\}$
The above order is the one introduced in [Tanemura93]*Section 4 with a relevant exception. Indeed in [Tanemura93] the author supposes that $(x_{1},x_{2},\dots,x_{n})$ is a path in ${\mathbb{Z}}^{2}$ but then refers to the ordering also when dealing with more general strings, as the above ones. If $(x_{1},x_{2},\dots,x_{n})$ is a path, then a(k)=k1a(k)=k-1 for all $k=1,2,\dots,n$ and the above definition of $\prec$ coincides with the order on $\Delta\{x_{1},x_{2},\dots,x_{n}\}$ introduced in [Tanemura93]*Section 4.
@@ 6.2 Exploration process
We denote $x_{1}^{s}=(0,s)$ for $s\in\{0,\dots,M-1\}$. Note that in the above notation $\Lambda_{L}^{\prime}$, $\partial_{l}\Lambda_{L}^{\prime}$ and $\partial_{r}\Lambda_{L}^{\prime}$ the integer $M$ is kept implicit. Suppose that on some probability space, with probability measure ${\mathbb{Q}}_{L}$ the following events are defined: $\{x_{1}^{s}:=(0,s)\text{ is occupied}\}$ for $s=0,1,\dots,M-1$, $\{y\text{ is linked to }x\}$ for |xy|=1|x-y|=1 and $x,y\in\Lambda^{\prime}_{L}$.
@@
[…]
@@
If $x_{1}^{s}$ is occupied (i.e. $E^{s}_{1}\not$$\emptyset$), then we proceed as follows. Suppose that we have defined all the sets preceding $C^{s}_{j+1}$ in the above string (34) (i.e. up to $C^{s}_{j}$), that we have not assigned to $J^{s}$ any value yet and that we want to define $C^{s}_{j+1}$. We call $W^{s}_{j}$ the points of $\Lambda_{L}^{\prime}$ involved in the construction up to this moment, i.e.
$W^{s}_{j}=\{x_{1}^{k}:0\leq k\leq M-1\}\cup\{x^{s^{\prime}}_{r}:0\leq s^{\prime}<s,\,1<r\leq\sharp\Lambda_{L}^{\prime}\}\cup\{x^{s}_{r}:1<r\leq j\}\,.$
We recall that it must be $E^{s}_{0}\subset E^{s}_{1}\subset\cdots\subset E^{s}_{j}$ and at each inclusion either the two sets are equal or the second one is obtained from the first one by adding exactly one point adjacent to the previous one. Hence, we can order the elements of $E^{s}_{j}$ according to the order with which the points have been added. This order on $E^{s}_{j}$ induces an order $\prec$ on the boundary $\Delta E^{s}_{j}$ as described in Section 6.1. We call $\mathcal{P}^{s}_{j}$ the following property: $E^{s}_{j}$ is disjoint from the right boundary $\partial_{r}\Lambda_{L}^{\prime}$ and $(\Lambda_{L}^{\prime}\cap\Delta{E}^{s}_{j})\setminus W^{s}_{j}\not$$\emptyset$. If property $\mathcal{P}^{s}_{j}$ is satisfied, then we define $x^{s}_{j+1}$ as the largest element of $(\Lambda_{L}^{\prime}\cap\Delta{E}^{s}_{j})\setminus W^{s}_{j}$ w.r.t. $\prec$. Let $k$ be the largest integer such that $x_{k}^{s}\in E^{s}_{j}$ and $|x^{s}_{j+1}-x^{s}_{k}|=1$. We decide where to add $x^{s}_{j+1}$ as follows: